OpenStudy (anonymous):

i've been asking for two hours, please help!! Which of the following is an equation for the horizontal asymptote to the graph of y=3-(x+b/x-c) and how you get it?

OpenStudy (anonymous):

the one that has a vertical asymptote at \(x=c\)

OpenStudy (anonymous):

oh you want "horizontal asymptote" right ?

OpenStudy (anonymous):

the answer is y=2, but i dont know how to solve for it.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

method one add

OpenStudy (anonymous):

drawing will be helpful, thank you

OpenStudy (anonymous):

\[3-\frac{x+b}{x-c}\] \[=\frac{3(x-c)-(x+b)}{x-c}\] \[=\frac{2x+stuff }{x-c}\]

OpenStudy (anonymous):

you can find the stuff, but no matter since the numerator has leading coefficient 2 and the denominator has leading coefficient 1 so the horizontal asymptote is \(y=2\)

OpenStudy (anonymous):

what do you do after step 3

OpenStudy (anonymous):

method 2 (no algebra, just think) \(-\frac{x+b}{x-c}\) has horizontal asymptote \(y=-1\) add \(3\) and get \(y=2\)

OpenStudy (anonymous):